OpenStudy (anonymous):

please .....Determine if the points (2,1) , (0,2) ,(4,0) are collinear.

OpenStudy (anonymous):

Equation of line passing through (0,2) and (4,0) Slope is -1/2 Line is y = -1/2 x + c 0 = -1/2 *4 + c 0 = -2 +c 2 = c Equation of line is y = -1/2 x + 2 Plug in x =2 and find y y = -1/2 * 2 + 2 = -1 +2 = 1 (2,1) lies on the line So the points (2,1) , (0,2) ,(4,0) are collinear.

OpenStudy (anonymous):

no it is'nt the value of the area covered is not zero

OpenStudy (anonymous):

Please check using a graph. All the points lie in one line so collinear

OpenStudy (anonymous):

will do thanks a lot